OpenStudy (anonymous):

(4(pir^3-600))/r^2=0 what is r?

OpenStudy (anonymous):

\[\frac{ 4(\pi r^3-600) }{ r^2 }=0\]

OpenStudy (anonymous):

\[\pi r ^{2} - 600 = 0\] \[r=\sqrt{600/\pi}\] \[r \approx 13.82\] or \[r \rightarrow -\infty\] or \[r \rightarrow \infty\]